7. The break -even in units sold will decrease if there is an increase in: b. Total fixed expenses e. Unit variable expenses d. Selling price e. None of the above Givens the following, the contributicn margin per unit is. , Sales (s6.000 unts) Variable Net operating income 8. the following, the contribution margin per unit is. $256,000 $160.000 a. $16 С.$2 prodacts on the basis of standard direct labor hours. If the labor 9. Variable manufacturing overhead is applied to variance is unfavorable, the variable overhead efficieney variance will be a Favorable b. Unfavorable c. Either favoeable or unfavorable d. zero 10. The Swenson Corp has the following data from their standard costing system: Actual quantity of DM purchased-35,000 pounds Standard price of DM- $4/pound Matcrial price variance $7,000 unfavorable Material quantity variance $4,200 favorable Then actual price per pound of direct materials purchased during the month is: a. $3.92 b. $4.32 c. $4.08 d. $4.20 11. A favorable labor rate variance indicates that a. Actual hours exceed standard hours b. Standard hours exceed actual hours e. The actual rate exeeeds the standard rate d. The standard rate exceeds the actual rate 12. A total of 6,850 ounces of a raw material was purchased at a total cost of $21,920. The materials price variance was $1,370, favorable. The standard price per ounce for the raw material must be a. $0.20 b. $0.90 С. $3.00 d. $3.40 e $3.20
Answer #1
ques 1
d. selling price
because the selling price will increase the contribution margin will increase
Breakeven point = Fixed cost / brekeven cost p.u
so your denominator increase so the breakeven comes down/decreases
ques 2
d.$6 per unit
         16,000
Sales $ 256,000.00 $        16.00
Variable expense $ 160,000.00 $        10.00
Contribution margin $    96,000.00 $          6.00
ques 3
b. unfavourable
because if efficiency variance is unfavourable means that
that actual hours are more than the standard hours for actual quantity
hence efficiency variance will be unfavourable
ques 4
d. $4.20
Materials price variance = AQ × (AP - SP)
$7,000 U = 35,000 pounds × (AP - $4 per pound)
$7,000 = 35,000 pounds × (AP - $4 per pound)
AP - $4 per pound = $7,000 ÷ 35,000 pounds
AP - $4 per pound = $0.20 per pound
AP = $4 per pound + $0.20 per pound
AP = $4.20 per pound
ques 5
d.the standard rate exceeds the actual rate
because the actual costs will be less than budgeted costs hence favourable
ques 6
d.$3.40
price variance = $1370 F
1370 = (AQ*SP)-(AQ*AP)
1370= (6850 *SP)-21920
23290/6850 K.GS = 3.40